OpenStudy (anonymous):

I don't know how to solve this. I am doing my homework ahead of time and I don't know ehat to do. Please help. Thanks! :D Write a linear function that satisfies the given conditions. 24. f(0) = 4, f(1) = 7

OpenStudy (anonymous):

a linear function is of the form f(x) = ax + b so we have f(0)= a*0 + b = b = 4 and f(1) = a*1 + b = a + 4 = 7 does that help?

OpenStudy (anonymous):

So what represents what? Thank you btw! :D

OpenStudy (anonymous):

so an example of a linear function would be f(x) = 5x +2 we have a number next to the x, which i call number "a" and a number added on as a constant term, number "b" f(x) = ax+b x is the input of the function, we put a number x in, and get another number f(x) out

OpenStudy (anonymous):

So how would I do that? Sorry, I'm just a little bit confused, I'm in 6th grade...

OpenStudy (anonymous):

So if I had this problem: g(–2) = 10 It would be: g(-2)=m(-2)+b=10

OpenStudy (anonymous):

When m=a

OpenStudy (anonymous):

yeah thats correct, a lot of people use m instead of a and c instead of b

OpenStudy (anonymous):

g(-2) = -2m + b

OpenStudy (anonymous):

Okay, thanks very much! :D

OpenStudy (anonymous):

no problem :)

OpenStudy (precal):